OpenStudy (anonymous):

Please factor the polynomial: 2x^2-6x+4

OpenStudy (uri):

2x^2-6x+4 2(x^2-3x+2)

OpenStudy (souvik):

=2{x^2-2x-x+2} =2{x(x-2)-(x-2)} =2(x-2)(x-1)